# An Efficient Wireless iBCI Headstage with Adaptive ADC Sample Rate

### 23 Apr 2026 v1

@@ V-B Compression efficiency
To assess the compression efficiency, we compare the compression ratio of our proposed adaptive sample rate mechanism with other two baselines in macaque and rat datasets. Fig. 7 indicates that our ADC adaptive sample system keeps the similar compression ratio to DCT and CS while decreasing the spike detection error rate by around 2%. Sample rate adjustment can not achieve comparable compression ratio to these two advanced compression algorithms, but electrode-adaptive sample rate allocation exhibits lower error than uniform compress the neural signal. Besides, adaptive spike detection threshold also provide more accurate spike activities in the same compression ratio compared with the uniform spike detection or noise-based detection.
@@ V-C Power efficiency
To evaluate the power efficiency of the proposed adaptive ADC sampling mechanism, we measured the power consumption during the streaming process using a power meter. Fig. 8 compares the power consumption across three datasets, demonstrating that our adaptive sampling method achieves the highest power efficiency. Specifically, the adaptive ADC sampling reduces power consumption by 24–35 mW compared to the DCT implementation on the FPGA layer. Furthermore, compared to CS, the adaptive approach yields even greater savings of 29–40 mW, as CS is computationally more intensive than DCT. Since the detected spike activities transmitted via Wi-Fi constitute a small data payload, the radio operates at a low duty cycle. Consequently, the majority of power is consumed by the Intan ADC chip and FPGA processing. Direct configuration of the ADC sampling rate proves to be more power-efficient because it primarily reduces the sampling operations of the ADC chip itself. Additionally, the FPGA layer only needs to process significant samples, resulting in reduced buffer usage and lower computational overhead.
@@ V-D Resource utilization efficiency
Figure 9: FPGA resource utilization percentage among three neural signal streaming schemes in rate dataset.
@@ V-E Impact on Neural Decoders
We further evaluated the downstream impact of our adaptive sampling wireless headstage on both visual and motor decoding tasks. The motor decoder translates neural spiking activity into the estimated trajectory of the mouse. As illustrated in Fig. 10, the proposed adaptive ADC sampling mechanism achieves the highest accuracy, with a median trajectory prediction error of 0.12 m. This represents a reduction in error of 0.02 m and 0.06 m compared to CS and DCT, respectively. The superior performance of our headstage is directly linked to its computational efficiency. While all methods maintain comparable compression ratios, the adaptive sampling approach imposes significantly lower signal processing overhead. In contrast, the high computational complexity of DCT and CS introduces processing latency, leading to buffer overflows and the loss of critical neural signal frames during high-activity periods. By minimizing these processing delays, our headstage scales effectively with data volume and MEA size, making it particularly suitable for closed-loop applications with strict latency requirements.
@@
Figure 10: Impact of adaptive sample, CS, and DCT on the motor decoder from the Cebra.
Figure 11: Impact of adaptive sample, CS, and DCT on the video decoder from the Cebra.
@@ VI Discussion and Limitations
Flexible sample rate modulation. A primary limitation of the current framework is that the sample rate modulation is restricted to integer downsampling factors (e.g., $1/2,1/3$). This coarse granularity prevents the system from precisely matching the target sampling rate to the signal’s sparsity. Consequently, the allocated sampling rate is often forced to the next available integer tier, leading to residual bandwidth inefficiency. Future iterations of this work will aim to implement fractional downsampling schemes (e.g., $3/4$) to achieve higher-resolution frequency adjustment and minimize over-sampling.
@@
Future Work: Acute In-Vivo recording. To bridge the gap between simulation and clinical application, we plan to transition from offline emulation to real-time in vivo experiments. In collaboration with the Department of Neuroscience at our institution, we will conduct acute implantation experiments on anesthetized mice. The choice of an acute, anesthetized model represents a strategic intermediate step before chronic implantation in behaving animals.
@@ VII Conclusion
In this work, we propose a wireless iBCI headstage featuring lightweight, adaptive control of ADC sampling rates and spike detection thresholds. Unlike conventional signal compression systems that rely on complex processing in the application layer, our approach directly reconfigures downsampling at the ADC layer. This server-driven design represents the first framework to compress data at the true source of the stream. Extensive experiments demonstrate that our headstage achieves significant savings in power and FPGA resource utilization while maintaining high compression efficiency and decoding accuracy.

### 24 Apr 2026 v2

@@ V-B Compression efficiency
To assess the compression efficiency, we compare the compression ratio of our proposed adaptive sample rate mechanism with the other two baselines on macaque and human datasets. As shown in Fig. 7, our ADC-level adaptive sampling achieves a compression ratio comparable to DCT and CS while reducing the spike detection error by approximately 2%. Although the absolute compression ratio of sample rate adjustment alone does not surpass these two transform-based algorithms, the electrode-adaptive allocation consistently yields lower detection error than uniformly compressing the neural signal. In addition, the adaptive spike detection threshold provides more accurate spike identification at the same compression ratio compared with uniform or noise-based thresholding.
@@ V-C Power efficiency
To evaluate the power efficiency of the proposed adaptive ADC sampling mechanism, we measured the power consumption during the streaming process using a power meter. Fig. 8 compares the power consumption across three datasets, demonstrating that our adaptive sampling method achieves the highest power efficiency. Specifically, the adaptive ADC sampling reduces power consumption by 24–35 mW compared to the DCT implementation on the FPGA layer. Furthermore, compared to CS, the adaptive approach yields even greater savings of 29–40 mW, as CS is computationally more intensive than DCT. Since the detected spike activities transmitted via Wi-Fi constitute a small data payload, the radio operates at a low duty cycle. Consequently, the majority of power is consumed by the Intan ADC chip and FPGA processing. Direct configuration of the ADC sample rate proves to be more power-efficient because it primarily reduces the sampling operations of the ADC chip itself. Additionally, the FPGA layer only needs to process significant samples, resulting in reduced buffer usage and lower computational overhead.
@@ V-D Resource utilization efficiency
Figure 9: FPGA resource utilization percentage among three neural signal streaming schemes in rat dataset.
@@ V-E Impact on Neural Decoders
We further evaluated the downstream impact of our adaptive sampling wireless headstage on both visual and motor decoding tasks. The motor decoder translates neural spiking activity into the estimated trajectory of the rat. As illustrated in Fig. 10, the proposed adaptive ADC sampling mechanism achieves the highest accuracy, with a median trajectory prediction error of 0.12 m. This represents a reduction in error of 0.02 m and 0.06 m compared to CS and DCT, respectively. The superior performance of our headstage is directly linked to its computational efficiency. While all methods maintain comparable compression ratios, the adaptive sampling approach imposes significantly lower signal processing overhead. In contrast, the higher computational complexity of DCT and CS may introduce processing latency that risks buffer overflows and the loss of neural signal frames during high-activity periods. By minimizing these processing delays, our headstage scales effectively with data volume and MEA size, making it particularly suitable for closed-loop applications with strict latency requirements.
@@
Figure 10: Impact of adaptive sample, CS, and DCT on the motor decoder from CEBRA.
Figure 11: Impact of adaptive sample, CS, and DCT on the video decoder from CEBRA.
@@ VI Discussion and Limitations
Flexible sample rate modulation. A primary limitation of the current framework is that the sample rate modulation is restricted to integer downsampling factors (e.g., $\times 2,\times 3$). This coarse granularity prevents the system from precisely matching the target sample rate to the signal’s sparsity. Consequently, the allocated sample rate is often forced to the next available integer tier, leading to residual bandwidth inefficiency. Future iterations of this work will aim to implement fractional downsampling schemes (e.g., $3/4$) to achieve higher-resolution frequency adjustment and minimize over-sampling.
@@
Server dependency and scalability. Our framework relies on a server to compute the per-electrode configuration. In practice, this optimization runs only during periodic recalibration intervals (on the order of minutes), not on every sample. Between recalibrations, the headstage operates autonomously using the last received schedule. If the wireless downlink is temporarily interrupted, the headstage continues streaming with the most recent configuration, gracefully degrading rather than failing. Regarding scalability, the per-electrode optimization is inherently independent across channels, allowing the server-side computation to parallelize linearly with channel count. The on-chip scheduling logic (modulo counters) also scales with negligible area overhead per electrode, making the architecture well-suited for next-generation MEAs with thousands of channels.
Future Work: Acute In-Vivo recording. To bridge the gap between simulation and clinical application, we plan to transition from offline emulation to real-time in vivo experiments. In collaboration with the Department of Neuroscience at our institution, we will conduct acute implantation experiments on anesthetized mice. The choice of an acute, anesthetized model represents a strategic intermediate step before chronic implantation in behaving animals. These experiments will validate three key aspects: (1) thermal safety of the headstage under continuous operation near cortical tissue, (2) real-time spike detection fidelity compared with the offline emulation results reported in this work, and (3) the stability of the adaptive configuration under physiological noise conditions that are absent in curated datasets.
@@ VII Conclusion
In this work, we propose a wireless iBCI headstage featuring lightweight, adaptive control of ADC sample rates and spike detection thresholds. Unlike conventional signal compression systems that rely on complex processing in the application layer, our approach directly reconfigures downsampling at the ADC layer. This server-driven design represents the first framework to compress data at the true source of the stream. Extensive experiments demonstrate that our headstage achieves significant savings in power and FPGA resource utilization while maintaining high compression efficiency and decoding accuracy.
VIII Ethics Statement
This study exclusively used publicly available datasets that were collected under protocols approved by the respective institutional review boards. No new experiments involving human subjects or animals are conducted as part of this work.